what is the unit of ryderbergs constant and how is the derivation of ryderburgs constant get.
 

Grade:11

1 Answers

jjijijijij
5 years ago
ryderberg’s constant has no unit it is a constant whose value is 1.1*10-15
it is calculated as;
we know that E=hc/lambda(wavelength)
and also energy of nth orbit =2.18*10-18 *Z2/N[z=atomic number and n=number of the orbit]’
by law of conservation of energy we can say that
 E=hc/lambda(wavelength)=2.18*10-18 *Z2/N2
1/wavelength=2.18*10-18 *Z2/N2/hc
1/wavelength=1.1*10-15 *Z2/N2